Find a formula for P = f(t), the size of the population that begins in year t = 0 with 2230 members and decreases at a 3.6 % annual rate. Assume that time is measured in years.
P = f(t) =

Answer by stanbon(75887)

Find a formula for P = f(t), the size of the population that begins in year t = 0 with 2230 members and decreases at a 3.6 % annual rate. Assume that time is measured in years.
P = f(t) = 2230(1-0.036)^t
f(t) = 2230*0.964^t
-----------------------------
Cheers,
Stan H.
================